OpenStudy (anonymous):

The leg of a right triangle is 3 units and the hypotenuse is 11 units. What is the length, in units, of the other leg of the triangle? 112 units square root 112 units 130 units square root 130 units

OpenStudy (anonymous):

@One098 @sleepyhead314

OpenStudy (sleepyhead314):

pythagorean theorem a^2 + b^2 = c^2 so a = root (c^2 - b^2) a = root ( (11)^2 - (3)^2 ) = ?

OpenStudy (anonymous):

a^2+b^2=c^2, a=3, c=11. so sqrt of 112=10.58

OpenStudy (anonymous):

so the answer would be (b)

OpenStudy (anonymous):

?

OpenStudy (sleepyhead314):

yes, B would be correct :)

OpenStudy (anonymous):

thank yall so much i wish i could give yall both medals